OpenStudy (anonymous):

Some Algebra 1 Questions.. Did I get these right? What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 3 7x - 5y = -11 is it "-2"? What is the value of the x variable in the solution to the following system of equations? 2x - y = 6 4x - 2y = 12 is it "x can be any number as there are infinitely many solutions to the system." or "0"? What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x - y = 3 Is the answer "3"? THANKS

OpenStudy (anonymous):

I figured someone could also check these for me? If you want, they should be easy enough. The graph of a system of equations will intersect at more than 1 point. Always The graph of a system of equations with the same slope will have no solutions. Never (or is it Sometimes...?) The graph of a system of equations will intersect at exactly 1 point. Sometimes

OpenStudy (ledah):

#1 is correct

OpenStudy (ledah):

I would assume if you did these the exact same way. you got it down. So its safe to say your correct

OpenStudy (anonymous):

Well, sometimes I do problems and I mess up something really simple or I use the same method and screw up, soooo it's not safe to say I'm correct XD

OpenStudy (ledah):

graph of a system. same thing applies if you know this well enough. its always best to go with your first answer is what my american government teacher said..

OpenStudy (anonymous):

I know that, but sometimes I just need to be checked. I get weird with math. o:

OpenStudy (ledah):

haha don't worry I'll see if somebody can check these @timo86m Please check these for her :o

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3x - 6y = 3 7x - 5y = -11 is it "-2"? yes but plug that back in to either equation to get the other :) x = -3, y = -2 good job :)

OpenStudy (anonymous):

oh wait you want just the y :D nvrmnd

OpenStudy (anonymous):

you are right good job :)

OpenStudy (anonymous):

On all of them? o.O

OpenStudy (anonymous):

2x - y = 6 4x - 2y = 12 x=x

OpenStudy (ledah):

Told you bugsey o_o

OpenStudy (anonymous):

doing one by one :)

OpenStudy (anonymous):

4x + 2y = 12 x - y = 3 x=3

OpenStudy (anonymous):

you are right on all of them :D

OpenStudy (anonymous):

sweet! :D Thank you!

OpenStudy (ledah):

told ya never doubt your first answer

OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6 @Sparklestaraa can you help

OpenStudy (triciaal):

this is such a long thread too many questions for one post!! parallel lines have the same slope no solution

OpenStudy (triciaal):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 3 7x - 5y = -11

OpenStudy (triciaal):